OpenStudy (anonymous):

MEDAL AND FAN PLEASE I need some help on an online school project. I honestly tried and I just dont understand I dont want awnsers just some help

OpenStudy (triciaal):

what do you need help with? do you have the question? if you post the question someone will be happy to help you

OpenStudy (anonymous):

@zepdrix

OpenStudy (triciaal):

part a think about y = mx + b y = function m = slope and b = intercept doing students each year 2ys ago then last year then this year change per year = 45 school 1 g(x)= 45x + 630 do the same for the other 2 schools

OpenStudy (triciaal):

part b find when x = 4, when x = 5 and when x = 6 for each function

OpenStudy (aum):

^^^ Be careful with part b). If you choose x = 0 to represent 2 years ago, then for part b, x should be 6, 7 and 8. But I'd prefer to choose x = 0 to represent the current year, x = -1 last year, x = -2 two years ago. Then g(x) = 45x + 720. If this equation is used, then for part b, x will be 4, 5 and 6.

OpenStudy (anonymous):

so far i have d(x)= 60x+630 h(x)=80x+600 for part a is this correct? @triciaal @aum

OpenStudy (triciaal):

yes part a

OpenStudy (triciaal):

@aum can you please finish this part c and d thanks.

OpenStudy (triciaal):

@Hero @ jim_thompson @zepdrix please help

OpenStudy (aum):

a) Let x represent the CURRENT year. g(x) = 45x + 720 d(x) = 60x + 750 h(x) = 80x + 760 b) In 4 years, g(4) = 45*4 + 720 = 900, d(4) = 60*4 + 750 = 990 h(4) = 80*4 + 760 = 1080 Complete the data when x = 5 and x = 6 c) Assuming Reya and Ian graduated this year and that they are attending a four year college: set x = -4 and calculate g(-4), d(-4), h(-4). Create a new column on the right to the given table, label the column, "Your Freshman Year", and enter below it g(-4), d(-4), h(-4). d) Look at the last column at the number of students that were in the Freshman class four years ago. Which one would your prefer based purely on the number of students that were in that class that year. State your reason. Smaller class size may mean more access to teacher? More students means better chances of finding compatible friends?